in solid mensuration find the length of the diagonal of (a) a square in terms of its edges a;(b) a rectangle in terms od its base b and its altitude h.
Answer by stanbon(75887)

in solid mensuration find the length of the diagonal of
(a) a square in terms of its edges a;
d = sqrt(a^2+a^2] = sqrt(2a^2) = sqrt(2)*a
==================================================
(b) a rectangle in terms od its base b and its altitude h.
d = sqrt(b^2 + h^2]
========================
Cheers,
Stan H.
========================